OpenStudy (anonymous):

Approximate the real number solutions to the polynomial function f(x) = x3 + 4x2 - 4x - 16

OpenStudy (boldjon):

f(x) = x^3 - 4x^2 -4x + 16 f(x) = x^2(x-4) -4(x-4) f(x) = (x-4)(x^2-4) 0 = (x-4)(x^2-4) x - 4 = 0 x^2 - 4 = 0 x = 4 x^2 = 4 x = ±2 Therefore, 3 solutions

OpenStudy (mathmale):

@boldjon: Remember: do NOT give away answers. Encourage the other person to find them himself. Re-read what the Code of Conduct says about that: http://openstudy.com/code-of-conduct

OpenStudy (boldjon):

oh.....i forgot *.*

OpenStudy (anonymous):

Thank you for explaining

OpenStudy (boldjon):

u welcoome